OpenStudy (anonymous):

I'm having problems understanding how to find the smaller segment of a circle. I know the formula but every time I try to solve the problem I'm not getting the correct answer. the chord is 7 and angle is 60 degrees

OpenStudy (dumbcow):



OpenStudy (dumbcow):

is what i drew what you are trying to solve for??

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

ok Area = area(sector) - area(triangle) we need to find radius  r = 7 because its an equilateral triangle area(sector) = pir^2 *(60/360) = 49pi/6 height of triangle = 7sqrt3/2 area(triangle) = 1/2 *7 *7sqrt3/2 = 49sqrt3/4 Area = 49pi/6 - 49sqrt3/4